Bramble Cole Inc. acquired the following assets in January of 2018.

Equipment, estimated service life, 5 years; salvage value, $14,800 $515,800
Building, estimated service life, 30 years; no salvage value $642,000


The equipment has been depreciated using the sum-of-the-years’-digits method for the first 3 years for financial reporting purposes. In 2021, the company decided to change the method of computing depreciation to the straight-line method for the equipment, but no change was made in the estimated service life or salvage value. It was also decided to change the total estimated service life of the building from 30 years to 40 years, with no change in the estimated salvage value. The building is depreciated on the straight-line method.

(a) Prepare the general journal entry to record depreciation expense for the equipment in 2021.
(b) Prepare the journal entry to record depreciation expense for the building in 2021.


(Round answers to 0 decimal places, e.g. 125. Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)

No.

Account Titles and Explanation

Debit

Credit

(a)

Enter an account title

Enter a debit amount

Enter a credit amount

Enter an account title

Enter a debit amount

Enter a credit amount

(b)

Enter an account title

Enter a debit amount

Enter a credit amount

Enter an account title

Enter a debit amount

Enter a credit amount
Answer #1

a) Accumulated depreciation Equipment = (515800-14800*12/15) = 400800

Depreciation expense for 2021 = (515800-400800-14800/2) = 50100

Accumulated depreciation for equipment = 642000/30*3 = 64200

Depreciation expense for 2021 = (642000-64200/37) = 15616

No.

Account Titles and Explanation

Debit

Credit

(a)

Depreciation expense 50100
Accumulated depreciation-Equipment 50100

(b)

Depreciation expense 15616
Accumulated depreciation-Building 15616